Could you please help me? On this problem, I come up with two answers: 2 and 2.5. Which is correct? 0.5/a=1.25/5. A=? Please advise. Thank you.
Answer by Alan3354(69443)

0.5/a=1.25/5
0.5/a = 0.25
0.5 = 0.25a
a = 0.5/0.25
a = 2
-----------
1 answer is sufficient